Let A be {1, 2, 3, 4}. Then the sequences 124, 421, 341 and 243 are same  permutations of A taken 3 at a time. The sequences 12, 43, 31, 24, and 21 are examples of different permutations of A taken two at a time.


1
Expert's answer
2021-12-20T16:42:06-0500

Solution:

Given, A ={1, 2, 3, 4}

Cyclic notations of given permutations are:

"(1\ 2\ 4)=\begin{pmatrix} 
1 &2&3&4
\\ 2& 4&3&1
 \end{pmatrix}"

"(4\ 2\ 1)=\begin{pmatrix} 
4 &2&1&3
\\ 2& 1&4&3
 \end{pmatrix}"

"(3\ 4\ 1)=\begin{pmatrix} 
3 &4&1&2
\\ 4& 1&3&2
 \end{pmatrix}"

"(2\ 4\ 3)=\begin{pmatrix} 
2 &4&3&1
\\ 4& 3&2&1
 \end{pmatrix}"

"(1\ 2)=\begin{pmatrix} 
1 &2&3&4
\\ 2& 1&3&4
 \end{pmatrix}"

"(4\ 3)=\begin{pmatrix} 
4 &3&1&2
\\ 3& 4&1&2
 \end{pmatrix}"

"(3\ 1)=\begin{pmatrix} 
3 &1&2&4
\\ 1& 3&2&4
 \end{pmatrix}"

"(2\ 4)=\begin{pmatrix} 
2 &4&3&1
\\ 4& 2&3&1
 \end{pmatrix}"

"(2\ 1)=\begin{pmatrix} 
2 &1&3&4
\\ 1& 2&3&4
 \end{pmatrix}"